The ages (in years) of the eight passengers on a bus are listed below.
9 1 29 11 22 46 40 35
Find the median age.
Select one:
A. 29 yr
B. 25.5 yr
C. 24.5 yr
D. 22 yr
Answer #1

Median is the middle term when the data is arranged in ascending order.

1, 9, 11, 22, 29, 35, 40, 46

There are 8 data values

Median = (n+1)/2 th value

= (8+1)/2 = 4.5th value

= average of 4th and 5th value

= (22 + 29)/2

= 25.5 yr